solve for r 3v=4(pi)r^3
Found 2 solutions by mananth, edjones:
Answer by mananth(16949)

3v=4(pi)r^3
divide by 4*pi
3v/4pi=r^3
r= cuberoot (3v/4*pi)

Answer by edjones(8007)

4(pi)r^3=3v
r^3=3v/4pi
r=(3v)^(1/3)/(2^(2/3)pi^(1/3))
=(6v)^(1/3)/2pi^(1/3)
.
Ed